OpenStudy (blenchiki):

(-i)^13 a. -13i b. -i c. -1 d. 1 e. i can someone explain how to solve? please and thank you

jhonyy9 (jhonyy9):

hope you know that i^2 = -1 yes ? so you can rewriting (-i)^13 = [(-i)^2]^6 *(-i) = [(-(-1))]^6 *(-i) = (1)^6 *(-i) = 1*(-i) = -i so choice b. is right sure hope helped

OpenStudy (blenchiki):

thank you @jhonyy9

jhonyy9 (jhonyy9):

was my pleasure hope you understand the way how you can solve the same exercises - yes ?

OpenStudy (blenchiki):

yes most definitely @jhonyy9

jhonyy9 (jhonyy9):

ok. great